# Mode analysis of Nambu-Goldstone modes in U(1) charged first-order relativistic viscous hydrodynamics

### v2

@@ 4.5 Transverse mode
The transverse mode is absent in a perfect fluid, as sound waves are “longitudinal” as we study in high school. However, in a viscous fluid, the situation is slightly different. Just like the longitudinal mode, the dispersion relation for the transverse mode is Hiscock:1985zz
$\omega(\rho+P)+i\kappa_{2}\omega^{2}+i\eta k^{2}=0.$ (252)
@@
$\omega_{1}=-i\Gamma_{\rm T}k^{2},~{}\Gamma_{\rm T}\equiv\frac{\eta}{\rho+P}.$ (253)
@@
Thus, $\kappa_{i}$ disappears in the first order, and the dispersion relation is hydrodynamical frame invariant. Since $\eta\geq 0$, ${\rm Im}~{}\omega<0$ if $\rho+P>0$. The transverse mode decouples from $\theta$, so the conclusion does not depend on whether charged or not. For $\rho+P=0$, we find $\kappa$ hydrodynamical frame independent. $\kappa\neq 0$ is a peculiar situation, and the transverse mode is unstable unless $\eta=0$. However, such a solution needs to be consistent with the perturbative expansion of $\omega$. Hence, the transverse mode does not exist for $\rho+P=0$. Thus, the null-energy condition is also a necessary condition for stability in the transverse mode.
4.6 Diffeomorphism invariance of Mode analysis
The critical benefit of our effective theory is that the diffeomorphism invariance of mode analysis is straightforward. Under a coordinate transformation $x^{\mu}\to x^{\prime\mu}=x^{\mu}+\xi^{\mu}(x)$, the spacetime NG modes transform as
$\pi^{\mu}(x)\to\pi^{\prime\mu}(x^{\prime})=\pi^{\mu}(x)+\xi^{\mu}(x).$ (254)
With this transformation, the fluid spacetime coordinates, i.e., the internal spacetime labels defining the fluid, are scalars in physical spacetime:
$\sigma^{\prime\alpha}(x^{\prime})=x^{\prime\alpha}-\pi^{\prime\alpha}(x^{\prime})=x^{\alpha}+\xi^{\alpha}(x)-(\pi^{\alpha}(x)+\xi^{\alpha}(x))=\sigma^{\alpha}(x).$ (255)
The local temperature is also a physical scalar, as $b$ in Eq. (164) is a scalar:
$b(x)=\sqrt{-\frac{\partial x^{\mu}}{\partial\sigma^{0}}\frac{\partial x^{\nu}}{\partial\sigma^{0}}g_{\mu\nu}(x)}=\sqrt{-\frac{\partial x^{\prime\mu}}{\partial\sigma^{\prime 0}}\frac{\partial x^{\prime\nu}}{\partial\sigma^{\prime 0}}g^{\prime}_{\mu\nu}(x^{\prime})}=1+\partial^{\prime}_{0}\pi^{\prime 0}(x^{\prime}).$ (256)
Thus, the coordinate transformations for the NG modes are nonlinear, and the action remains invariant under these transformations by combining terms of different orders in $\partial\pi$. Similarly, the fluid velocity transforms as follows:
$u^{\mu}(x)\to u^{\prime\mu}(x^{\prime})=\frac{\partial x^{\prime\mu}}{\partial x^{\nu}}u^{\nu}(x)=\frac{\partial x^{\prime\mu}}{\partial x^{\nu}}\frac{\partial x^{\nu}}{\partial\sigma^{0}}b^{-1}=\frac{\partial x^{\prime\mu}}{\partial\sigma^{\prime 0}(x^{\prime})}b^{-1}=(1,\partial^{\prime}_{0}\pi^{\prime i}(x^{\prime})).$ (257)
Covariant derivatives, external fields, and the noise NG modes transform linearly. The effective action (160) is constructed from these covariant quantities.
Now, let us consider the Lorentz boost transformation as a special case of $\xi^{\mu}(x)=\Lambda^{\mu}{}_{\nu}x^{\nu}-x^{\mu}$. Under this transformation, the Minkowski metric remains unchanged, so the effective action for the NG modes does not change. Thus, the mode analysis in the transformed coordinates is equivalent to the original one.
In a general curved background, one may take a local Minkowski frame, and the mode analysis remains the same with the present case within the curvature radius, i.e., for $\omega,k\gg\sqrt{R}$, where $R$ is the Ricci curvature. For $\omega,k<\sqrt{R}$, one must use a proper cosmological background, where the dynamics of the NG modes depend on the geometric structure.
This discussion implies that the sound speed $c_{s/n}$ is unchanged under diffeomorphism transformations. The sound speed is defined relative to the rest frame of fluids, which should be independent of the choice of frame for the mode analysis. This is consistent with the standard notion of sound waves, where Lorentz transformations induce the Doppler shift. The sound speed measured by a non-comoving observer is distinct from the intrinsic sound speed defined with respect to the fluid rest frames, as the former includes a Doppler effect. In other words, the relative sound velocity is incompatible with the present covariant theory for hydrodynamical modes.
@@ 5.1 Green functions
One finds the reduced noise matrix (217) by eliminating the redundant freedom. The matrix components in Eq. (222) are found as
$\tilde{\mathcal{N}}^{\rm L}=\frac{ik^{2}(\rho+P)\Gamma_{\rm L}}{2\beta}+\mathcal{O}(\varepsilon^{2}),$ (258)
@@
$\tilde{\mathcal{N}}^{\rm T}$ $=\frac{ik^{2}(\rho+P)\Gamma_{\rm T}}{4\beta}+\mathcal{O}(\varepsilon^{2}),$ (259)
@@
~$\tilde{\mathcal{E}}^{\rm L}$ $=\frac{1}{2}(\rho+P)(\omega-c_{s/n}k+i\Gamma_{\rm L}k^{2})(\omega+c_{s/n}k+i\Gamma_{\rm L}k^{2})+\mathcal{O}(\varepsilon^{2}),$ (260)
~$\tilde{\mathcal{E}}^{\rm T}$ $=\frac{1}{2}(\rho+P)(\omega+i\Gamma k^{2})+\mathcal{O}(\varepsilon^{2}).$ (261)
@@
$\frac{\tilde{\mathcal{N}}}{\tilde{\mathcal{E}}\tilde{\mathcal{E}}^{*}}=\frac{2}{\beta\omega}\left(\frac{1}{\tilde{\mathcal{E}}^{*}}-\frac{1}{\tilde{\mathcal{E}}}\right).$ (262)
Thus, the KMS condition for the full action guarantees the KMS condition for the dynamical fields after solving the constraint equations. Eq. (262) is understood as another representation of the dynamical KMS condition proposed in Ref. Glorioso:2017fpd . $\tilde{G}^{-1}$ is invertible and we find
$G=\begin{pmatrix}-\tilde{\mathcal{E}}^{{\rm T}-1}\tilde{\mathcal{N}}^{\rm L}\tilde{\mathcal{E}}^{{\rm L}*-1}&0&\tilde{\mathcal{E}}^{{\rm L}-1}&0\\ 0&-\tilde{\mathcal{E}}^{{\rm L}-1}\tilde{\mathcal{N}}^{\rm L}\tilde{\mathcal{E}}^{{\rm L}*-1}&0&\tilde{\mathcal{E}}^{{\rm T}-1}\\ \tilde{\mathcal{E}}^{{\rm T}*-1}&0&0&0\\ 0&\tilde{\mathcal{E}}^{{\rm L}*-1}&0&0\end{pmatrix}\equiv\begin{pmatrix}G_{K}&G_{R}\\ G_{A}&0\end{pmatrix}.$ (263)
$G_{K}$, $G_{R}$ and $G_{A}$ are the Keldysh (statistical), retarded and advanced propagators, respectively. The effective action for the NG mode coincides with the MSR action for the Brownian motion (299). By taking
$m\to\frac{\rho+P}{2},~{}\omega_{0}^{2}=c_{s/n}^{2}k^{2},~{}\gamma\to 2k^{2}\Gamma_{\rm L},$ (264)
in Eqs. (297) and (298), we find
$G_{R}(t,k)$ $=\Theta(t)\frac{e^{-k^{2}\Gamma_{\rm L}t}}{c_{s/n}k}\sin\left(c_{s/n}kt\right),$ (265)
$G_{K}(t,k)$ $=\frac{2e^{-k^{2}\Gamma_{\rm L}t}}{\beta(\rho+P)c_{s/n}^{2}k^{2}}\left(\cos\left(c_{s/n}kt\right)+\frac{\Gamma_{\rm L}k}{c_{s/n}}\sin\left(c_{s/n}kt\right)\right).$ (266)
Note that the dispersion relations are reliable only up to $\mathcal{O}(\varepsilon)$. The retarded propagator is constrained to $t>0$. Combining these Eqs. (265) and (266) with Eqs. (172) and (173), one can compute any correlation functions of $T$ and $J$.
These Green functions are identical to those found in Brownian motion bounded in a harmonic potential, with similar expressions derived in the context of hydrodynamics Kovtun:2012rj . However, this work explicitly confirmed the hydrodynamical frame invariance of these Green functions.
@@ 5.2 Comments on causal structure
Superluminal modes may arise when higher-order derivatives are included inconsistently. However, even when the effective theory is properly constructed, there are two key aspects of causality to address.
First, consider the causal structure with respect to the sound speed. For nontrivial dispersion relations, there is a subtlety in defining the propagation speed of waves and signals, such as phase velocity, group velocity, front velocity, etc. In particular, the standard definition of the front velocity is incompatible with hydrodynamics as it breaks down in the $k\to\infty$ limit. In the present formulation, we have the retarded Green’s function for stochastic NG modes, which explicitly defines the correlations between two distant spacetime points.
We first discuss the causal structure for $\Gamma_{\rm L}=0$. In this case, the retarded Green’s function (265) is equivalent to that of a relativistic scalar field in Minkowski spacetime, with $c=c_{s/n}$. The retarded propagator must vanish for spacelike separations with respect to $c_{s/n}$. In other words, we define $r_{s/n}\equiv c_{s/n}t-x$, and $G_{R}$ vanishes when $r_{s/n}<0$, which defines the sound cone. This reinforces the fact that sound waves cannot propagate faster than the sound speed, ensuring causality with respect to the sound speed.
Secondly, consider the magnitude of the adiabatic sound speed. Thermodynamic variables determine $c_{s/n}$, which remains unconstrained. While fast modes cannot travel faster than the speed of light, the superluminal propagation of sound waves is not forbidden by the speed limit of fast modes, as it does not involve the transfer of the fast modes themselves.
The stability condition $\Gamma_{\rm L}\geq 0$ ensures that $G_{R}(t)=0$ for $t<0$ when performing the inverse Fourier transform with respect to $\omega$. This is a necessary condition for the causal propagation of sound waves. Mathematically, this condition holds when the zeros of ~$\tilde{\mathcal{E}}^{\rm L}$ do not extend into the region ${\rm Im}~{}\omega>0$. For real $k$, we find
${\rm Im}~{}\omega=-\Gamma_{\rm L}k^{2}<0.$ (267)
The authors of Ref. Heller:2022ejw extended this inequality for complex wavenumbers, showing ${\rm Im}~{}\omega(k)\leq|{\rm Im}~{}k|$ by inspecting the analytic structure of a stable retarded Green’s function. Generalizing Eq. (267), they derived
${\rm Im}~{}\left[\pm c_{s/n}k-i\Gamma_{\rm L}k^{2}\right]<|{\rm Im}~{}k|.$ (268)
For $k=ip$ with $p\in\Re$ and $p\to 0$, we find $c_{s/n}\leq 1$. Since $\Gamma\geq 0$ is guaranteed by unitarity, the KMS condition, and the null energy condition, causality in first-order hydrodynamics is upheld by these principles. The relationship between causality and stability is also discussed in Refs. Olson:1990rzl ; Bemfica:2020zjp ; Heller:2022ejw ; Gavassino:2023myj .
Next, let us consider $\Gamma_{\rm L}>0$. In this case, the sound wave is dissipative, and the concept of causality becomes subtle. One can further Fourier transform the Green functions and find the Green functions in real space:
@@
$=\frac{\Theta(t)}{(4\pi)^{\frac{3}{2}}c_{s/n}|\mathbf{x}-\mathbf{y}|\sqrt{\Gamma_{\rm L}t}}e^{-\frac{(c_{s/n}t-|\mathbf{x}-\mathbf{y}|)^{2}}{4\Gamma_{\rm L}t}}.$ (269)
@@
In the $\Gamma_{\rm L}\to 0$ limit, the Gaussian distribution converges to the delta function, and the sound waves are constrained on the sound cone. However, nonvanishing $\Gamma_{\rm L}$ implies that the retarded propagator is nonzero for $r_{s/n}<0$, so that the correlation function extends outside of the sound cone. This is the causality problem in first-order hydrodynamics. The Gaussian distribution has non-vanishing tails for $|\mathbf{x}-\mathbf{y}|\to\infty$, which may naively lie outside the light cone. This is manifest even for the transverse mode with zero sound velocity, hence the bound on $c_{s/n}$ is insufficient to resolve this issue. However, it is evident that there is a validity range in the effective description (269) since hydrodynamics is an effective theory.
The width of the wave packet grows as $\sqrt{\Gamma_{\rm L}t}$, which corresponds to the MFP of the fast mode. The diffusion occurs due to the propagation of fast modes rather than sound waves as the net mass is transported. In fact, the Gaussian wave packet in Eq. (269) implies that the fast modes’s typical traveling distance is given by the MFP in the rest frame of sound waves. The maximum speed of the fast mode is constrained in the UV theory, which should not be beyond the light speed. Hence, one cannot naively extend Eq. (269) to $|\mathbf{x}-\mathbf{y}|\to\infty$. The center of the sound packet aligns with the sound cone. In the rest frame of sound waves, the fast modes diffuse with a maximum speed $c_{\rm fast}\leq 1$. Thus, the maximum front speed $c_{\rm front}$ of the diffused sound wave is given by the composition of $c_{s/n}\leq 1$ and $c_{\rm fast}\leq 1$, which is not the sum $c_{\rm front}\neq c_{\rm fast}+c_{s/n}$ but
$c_{\rm front}=\frac{c_{\rm fast}+c_{s/n}}{1+c_{\rm fast}c_{s/n}}\leq 1,$ (270)
@@
as we know in special relativity. The front speed never exceeds the light speed, thereby preserving causality in terms of the light cone, while that for the sound cone might be violated.
6 Conclusion
This paper investigated the dynamics of collective excitations around relativistic hydrodynamical solutions, incorporating noise effects, based on the Schwinger-Keldysh path integral formalism for a UV theory with global $U(1)$ symmetry. We employed the effective field theory (EFT) of dissipative fluids developed in Refs. Crossley:2015evo ; Glorioso:2017fpd , particularly the physical spacetime formulation at the semiclassical order.
In this framework, the collective modes are regarded as Nambu-Goldstone (NG) modes associated with the spontaneous symmetry breaking of global $U(1)$ and diffeomorphism symmetries. We derived a quadratic action for the NG modes in $U(1)$-charged, relativistic, dissipative hydrodynamics. The EFT coefficients are constrained by various conditions imposed on the path integral of the UV model. For example, the local KMS condition and the unitarity of the generating functional lead to the non-negativity of the transport coefficients. We derived the dispersion relation for the NG modes by eliminating redundant modes, solving the constraint equations, and discussing their stability and causality structure.

### v3

@@ 5.5 Transverse mode
The transverse mode is absent in a perfect fluid, as sound waves are “longitudinal” as we study in high school. However, in a viscous fluid, the situation is slightly different. Just like the longitudinal mode, the dispersion relation for the transverse mode is [11]
$\omega(\rho+P)+i\kappa_{2}\omega^{2}+i\eta k^{2}=0.$ (5.71)
@@
$\omega_{1}=-i\Gamma_{\rm T}k^{2},~\Gamma_{\rm T}\equiv\frac{\eta}{\rho+P}.$ (5.72)
@@
Thus, $\kappa_{i}$ disappears in the first order, and the dispersion relation is hydrodynamical frame invariant. Since $\eta\geq 0$, ${\rm Im}~\omega<0$ if $\rho+P>0$. The transverse mode decouples from $\theta$, so the conclusion does not depend on whether charged or not. For $\rho+P=0$, we find $\kappa$ hydrodynamical frame independent. $\kappa\neq 0$ is a peculiar situation, and the transverse mode is unstable unless $\eta=0$. However, such a solution needs to be consistent with the perturbative expansion of $\omega$. Hence, the transverse mode does not exist for $\rho+P=0$. Thus, the $\rho+P\geq 0$ is also a necessary condition for stability in the transverse mode.
5.6 Mode analysis in non-comoving inertial frames
So far we assumed a specific inertial frame where the fluid is comoving. The original UV theory enjoys the full Poincaré symmetry in flat spacetime, including Lorentz boosts. The presence of $\beta^{\mu}$ in the ensemble spontaneously breaks this symmetry. Then, a natural question is if the mode analysis in a non-comoving inertial frame is equivalent to that in the rest frame.
The answer to this question is straightforward as we employed the nonlinear realization theory. The broken boost symmetry is nonlinearly realized by the NG field $\pi^{\mu}$ through Eq. (4.6). Therefore, the dynamical theory of hydrodynamical modes are manifestly covariant under the nonlinear Lorentz boost.
To illustrate this more clearly, we consider an infinitesimal Lorentz boost in the $\tau$$x$ plane, generated by a tangent vector $\xi^{\mu}=v(x\delta^{\mu}_{0}+\tau\delta^{\mu}_{1})$. Under this transformation, the Minkowski background remains invariant, while the NG fields transform as $\pi^{0}\to\tilde{\pi}^{0}=\pi^{0}+vx$ and $\pi^{1}\to\tilde{\pi}^{1}=\pi^{1}+v\tau$. Although $\dot{\tilde{\pi}}^{0}=\dot{\pi}^{0}$, the time derivative of the spatial component changes: $\dot{\tilde{\pi}}^{1}\neq\dot{\pi}^{1}$. Consequently, the effective action in the boosted frame is invariant when expressed in terms of $\tilde{\pi}^{\mu}$.
This nonlinear realization implies that the dynamics of sound waves in a fluid at rest are physically equivalent to those in a boosted frame. More precisely, after the boost transformation, one identifies the fluid comoving with the new inertial frame, and hydrodynamical modes defined in that fluid is physically equivalent to the original modes. In this sense, the theory of hydrodynamical modes is always defined in a comoving inertial frame. This contrasts with a naive application of Lorentz boosts to the wavevector $k_{\mu}$. While an identification by the boost is intuitive, it is only physically meaningful when the field variables are linear representations of Lorentz boosts.
The theory formulated using $\pi^{\mu}$ directly in the boosted frame was analyzed in Ref. [11] in a traditional perturbation theory, and then instabilities were reported. It is now clear that this formulation is not dynamically equivalent to that in the rest frame, and the emergence of instabilities is not surprising. The dynamically equivalent theory in the boosted frame is instead formulated in terms of $\tilde{\pi}^{\mu}$, whose mode structure is manifestly identical to that in the original rest frame. The theory of $\pi^{\mu}$ in the non-comoving frame mixes density waves with mass transfer due to the moving frame, leading to a misidentification of the dynamical degrees of freedom.
Section summary.
This section presents the main contribution of this paper. Building on the effective action derived in the previous section, we performed a mode analysis of the hydrodynamical excitations. We carefully eliminated the redundancy in the NG modes by solving the constraint equations up to noise order, where noise fields correspond to the advanced-sector counterparts in the Schwinger-Keldysh formalism. The analysis was carried out within a consistent derivative expansion scheme. We emphasized that the derivative expansion of the constitutive relations is not, in general, aligned with that of the NG mode dynamics under generic hydrodynamical frames. A consistent truncation of derivative terms yields a manifestly frame-invariant mode structure, which, for linear modes, coincides with the Landau frame result. Upon imposing the KMS and unitarity conditions, we demonstrated that all hydrodynamical modes remain stable under the physical condition $\rho+P>0$.
@@ 6 Green functions and causality
This is the benefit of our formulation. We have derived the noise action in a general hydrodynamic frame, and carefully eliminate the redundant variables by solving the constraint equations. Consequently, we have the noise action written by the hydrodynamical frame invariant. To be more specific, the linear analysis can determine the advanced and retarded propagators, $G_{A}$ and $G_{R}$. In previous works, one may find the Keldysh propagator $G_{K}$ by the KMS condition. In our framework, we can find $G_{K}$ from the second order action, and the KMS condition among the propagators are derived as discussed shortly.
@@ 6.1 Green functions
One finds the reduced noise matrix (5.39) by eliminating the redundant freedom. The matrix components in Eq. (5.43) are found as
$\tilde{\mathcal{N}}^{\rm L}=\frac{ik^{2}(\rho+P)\Gamma_{\rm L}}{2\beta}+\mathcal{O}(\varepsilon^{2}),$ (6.1)
@@
$\tilde{\mathcal{N}}^{\rm T}$ $=\frac{ik^{2}(\rho+P)\Gamma_{\rm T}}{4\beta}+\mathcal{O}(\varepsilon^{2}),$ (6.2)
@@
~$\tilde{\mathcal{E}}^{\rm L}$ $=\frac{1}{2}(\rho+P)(\omega-c_{s/n}k+i\Gamma_{\rm L}k^{2})(\omega+c_{s/n}k+i\Gamma_{\rm L}k^{2})+\mathcal{O}(\varepsilon^{2}),$ (6.3)
~$\tilde{\mathcal{E}}^{\rm T}$ $=\frac{1}{2}(\rho+P)(\omega+i\Gamma k^{2})+\mathcal{O}(\varepsilon^{2}).$ (6.4)
@@
$\frac{\tilde{\mathcal{N}}}{\tilde{\mathcal{E}}\tilde{\mathcal{E}}^{*}}=\frac{2}{\beta\omega}\left(\frac{1}{\tilde{\mathcal{E}}^{*}}-\frac{1}{\tilde{\mathcal{E}}}\right).$ (6.5)
Thus, the KMS condition for the full action guarantees the KMS condition for the dynamical fields after solving the constraint equations. $\tilde{G}^{-1}$ is invertible and we find
$G=\begin{pmatrix}-\tilde{\mathcal{E}}^{{\rm T}-1}\tilde{\mathcal{N}}^{\rm L}\tilde{\mathcal{E}}^{{\rm L}*-1}&0&\tilde{\mathcal{E}}^{{\rm L}-1}&0\\ 0&-\tilde{\mathcal{E}}^{{\rm L}-1}\tilde{\mathcal{N}}^{\rm L}\tilde{\mathcal{E}}^{{\rm L}*-1}&0&\tilde{\mathcal{E}}^{{\rm T}-1}\\ \tilde{\mathcal{E}}^{{\rm T}*-1}&0&0&0\\ 0&\tilde{\mathcal{E}}^{{\rm L}*-1}&0&0\end{pmatrix}\equiv\begin{pmatrix}G_{K}&G_{R}\\ G_{A}&0\end{pmatrix}.$ (6.6)
$G_{K}$, $G_{R}$ and $G_{A}$ are the Keldysh (statistical), retarded and advanced propagators, respectively. The effective action for the NG mode coincides with the MSR action for the Brownian motion (C.11). By taking
$m\to\frac{\rho+P}{2},~\omega_{0}^{2}=c_{s/n}^{2}k^{2},~\gamma\to 2k^{2}\Gamma_{\rm L},$ (6.7)
in Eqs. (C.9) and (C.10), we find
$G_{R}(t,k)$ $=\Theta(t)\frac{e^{-k^{2}\Gamma_{\rm L}t}}{c_{s/n}k}\sin\left(c_{s/n}kt\right),$ (6.8)
$G_{K}(t,k)$ $=\frac{2e^{-k^{2}\Gamma_{\rm L}t}}{\beta(\rho+P)c_{s/n}^{2}k^{2}}\left(\cos\left(c_{s/n}kt\right)+\frac{\Gamma_{\rm L}k}{c_{s/n}}\sin\left(c_{s/n}kt\right)\right).$ (6.9)
Note that the dispersion relations are reliable only up to $\mathcal{O}(\varepsilon)$. The retarded propagator is constrained to $t>0$. Combining these Eqs. (6.8) and (6.9) with Eqs. (4.29) and (4.30), one can compute any correlation functions of $T$ and $J$.
These Green functions are identical to those found in Brownian motion bounded in a harmonic potential, with similar expressions derived in the context of hydrodynamics [1]. However, this work explicitly confirmed the hydrodynamical frame invariance of these Green functions.
@@ 6.2 Comments on causal structure
Superluminal modes may arise when higher-order derivatives are introduced inconsistently. However, even when the effective theory is properly constructed, there remain two key aspects of causality to address.
We first consider the constraint on the adiabatic sound speed. Thermodynamic variables determine $c_{s/n}$, which is, in principle, unconstrained. While fast modes in theories with Lorentz boost symmetry cannot propagate faster than the speed of light, the superluminal propagation of sound waves is not necessarily forbidden, as it does not involve the transfer of fast modes themselves.
The stability condition $\Gamma_{\rm L}\geq 0$ ensures that $G_{R}(t)=0$ for $t<0$ when performing the inverse Fourier transform with respect to $\omega$. Mathematically, this condition is satisfied when the zeros of ~$\tilde{\mathcal{E}}^{\rm L}$ do not extend into the region ${\rm Im}~\omega>0$. For real $k$, we find
${\rm Im}~\omega=-\Gamma_{\rm L}k^{2}<0.$ (6.10)
The authors of Ref. [32] extended this inequality to complex wavenumbers, showing that ${\rm Im}~\omega(k)\leq|{\rm Im}~k|$ by analyzing the analytic structure of a stable retarded Green’s function. Generalizing Eq. (6.10), they derived
${\rm Im}~\left[\pm c_{s/n}k-i\Gamma_{\rm L}k^{2}\right]<|{\rm Im}~k|.$ (6.11)
For $k=ip$ with $p\in\mathbb{R}$ and $p\to 0$, this yields $c_{s/n}\leq 1$. Since $\Gamma_{\rm L}\geq 0$ is guaranteed by unitarity, the KMS condition, and the null energy condition, the upper bound on the adiabatic sound speed in first-order hydrodynamics is supported by these principles. The connection between causality and stability is further discussed in Refs. [33, 34, 32, 35].
Next, we consider the causal structure with respect to the sound speed. For nontrivial dispersion relations, there is a subtlety in defining the propagation speed of waves and signals, such as the phase velocity, group velocity, front velocity, etc. In particular, the standard definition of front velocity is incompatible with hydrodynamics, as it breaks down in the $k\to\infty$ limit.
For $\Gamma_{\rm L}=0$, the retarded Green’s function (6.8) reduces to that of a relativistic scalar field in Minkowski spacetime with $c=c_{s/n}$. The retarded propagator must vanish for spacelike separations with respect to $c_{s/n}$. In other words, defining $r_{s/n}\equiv c_{s/n}t-x$, we find that $G_{R}$ vanishes for $r_{s/n}<0$, which defines the sound cone. This confirms that sound waves cannot propagate faster than the sound speed, ensuring causality with respect to the sound cone.
For $\Gamma_{\rm L}>0$, the sound wave becomes dissipative, and the notion of causality becomes more subtle. One can further Fourier transform the Green’s functions in $k$:
@@
$=\frac{\Theta(t)}{(4\pi)^{3/2}c_{s/n}|\mathbf{x}-\mathbf{y}|\sqrt{\Gamma_{\rm L}t}}e^{-\frac{(c_{s/n}t-|\mathbf{x}-\mathbf{y}|)^{2}}{4\Gamma_{\rm L}t}}.$ (6.12)
@@
In the $\Gamma_{\rm L}\to 0$ limit, the Gaussian distribution approaches a delta function, and the sound waves are constrained to the sound cone. However, a nonvanishing $\Gamma_{\rm L}$ implies that the retarded propagator becomes nonzero for $r_{s/n}<0$, meaning the correlation function extends outside the sound cone. This is the well-known causality issue in first-order hydrodynamics. The Gaussian distribution has nonzero tails for $|\mathbf{x}-\mathbf{y}|\to\infty$, which may naively extend beyond the light cone. This behavior is evident even for transverse modes with zero sound speed, indicating that bounding $c_{s/n}$ is insufficient to resolve the issue.
However, it is evident that Eq. (6.12) has a limited validity range, as hydrodynamics is an effective theory based on the derivative expansion. In Eq. (6.12), the integration is formally performed from $k=0$ to $\infty$ to obtain the real-space expression, but the exponential damping is only valid at leading order and does not hold in the $k\to\infty$ limit. Thus, the superluminal Gaussian tail does not imply a physical violation of causality but rather reflects the breakdown of the effective theory at short distances.
Therefore, the support of the retarded propagator in real space cannot be interpreted literally as a measure of causality. The width of the wave packet grows as $\sqrt{\Gamma_{\rm L}t}$, corresponding to the MFP of the fast modes. Diffusion occurs due to the propagation of these fast modes, not due to sound waves, as they carry net mass. In fact, the Gaussian wave packet in Eq. (6.12) suggests that the typical travel distance of fast modes is set by the MFP in the rest frame of the sound wave.
The maximum speed of the fast modes is determined by the UV theory and must not exceed the speed of light. Thus, Eq. (6.12) cannot be naively extrapolated to $|\mathbf{x}-\mathbf{y}|\to\infty$. The center of the sound packet still travels along the sound cone. In the rest frame of the sound wave, the fast modes diffuse with a maximum speed $c_{\rm fast}\leq 1$. Therefore, the maximum front speed $c_{\rm front}$ of the diffused sound wave is given by the relativistic addition of $c_{s/n}$ and $c_{\rm fast}$:
$c_{\rm front}=\frac{c_{\rm fast}+c_{s/n}}{1+c_{\rm fast}c_{s/n}}\leq 1,$ (6.13)
@@
as dictated by special relativity. The front speed never exceeds the speed of light, thereby preserving causality with respect to the light cone, even though causality with respect to the sound cone may be violated.
Section summary.
This section focused on the analysis at noise order in the effective action. We derived the Keldysh Green function from the effective action constructed in the previous sections. By consistently eliminating the redundancy of the NG modes up to quadratic order in the noise fields, the resulting noise action remains frame invariant. Upon imposing the KMS condition at the level of the effective action, we verified that the Green functions satisfy the standard KMS relation. We also examined the causal structure of the theory with respect to the sound cone. In particular, we discussed that the adiabatic sound speed is bounded by the speed of light. Furthermore, the front velocity of dissipative processes—defined as the relativistic composition of the rest-frame velocity and the mass transport velocity in that frame—is also constrained by the causal structure, ensuring that all signal propagation remains subluminal in hydrodynamics.
7 Conclusions
We have formulated a symmetry-based effective theory for relativistic hydrodynamics with a global $U(1)$ symmetry, valid up to semi-classical noise order. Our construction relies solely on the input of a local four-vector $\beta^{\mu}$, without assuming any thermodynamic relations at the outset. By imposing local KMS and unitarity conditions, we have systematically derived thermodynamic relations and fluctuation–dissipation constraints directly from the effective action.
Although the foundational structure of this construction was already presented in the seminal work [15], a key contribution of this paper is the identification of the quadratic action for hydrodynamical modes in the most general $U(1)$ hydrodynamics. The hydrodynamical mode arises as a nonlinear realization of spacetime symmetries, which makes the covariance of perturbation theory manifest. Our treatment also consistently incorporates noise contributions.
We performed a mode analysis by carefully eliminating unphysical redundancies. A consistent truncation of derivative terms yields a manifestly hydrodynamic-frame-invariant mode structure, which reproduces the Landau frame result in the linear regime. Upon imposing the KMS and unitarity conditions, we demonstrated that all hydrodynamical modes remain stable under the physical condition $\rho+P>0$. The dominant energy condition, or the positivity of the enthalpy density is the minimal requirement for the stability of first-order hydrodynamics in flat spacetime.